Solve for A in B = 2 divided by A-10

Answer by checkley71(8403)

if your answers are: 2/B-10 or 2/B + 10 or [2/B - 10] or 10 - 2/B
then:
B=2/(A-10) CROSS MULTIPLY
B(A-10)=2
AB-10B=2
B(A-10)=2
(A-10)=2/B
A=2/B-10